# Exploiting Negative Capacitance for Unconventional Coulomb Engineering

### v1

@@ IV Results
In addition to $|\eta_{f}|L_{f}$, two important length scales in the composite structure are $2k_{F}$ (low-energy scattering shell in the 2DES) and a characteristic length scale corresponding to the NC phenomenon, taken to be the domain width $d$ of PDT in the extrinsic case. We define two distinct regimes based on the relative values of these length scales: long-wavelength and short-wavelength.
In the long-wavelength limit, two simplifications emerge: wavelengths of electron scattering in 2DES are too large to resolve microscopic details on the scale of domain width $d$ (so that a simplified model with constant $\varepsilon^{nc}(q)$ suffices), and gate screening reduces to its linear limit since the gates see an effectively uniform charge distribution in the 2DES. Expressed as mathematical conditions:
  • In the long-wavelength regime, $2k_{F}d\ll 1$ and $2k_{F}|\eta_{f}|L_{f}\ll 1$, so that the single-mode approximation discussed in previous section is a reasonable starting point, and the simplified description of Eq. (3) is applicable [see Fig. 2(b)].
  • In the short-wavelength regime, $2k_{F}d\gg 1$ and $2k_{F}|\eta_{f}|L_{f}\gg 1$ , indicating that it is important to model dispersion $\varepsilon^{f}_{z}(q)$ (single-mode approximation may not necessarily be reliable) and gate screening modifies the $q$-space structure of $V_{\text{eff}}$ significantly [see Fig. 2(c)].
Figs. 2(b),(c) demonstrate how the two length scale regimes are accessed using 2DES carrier density $n$ as the control parameter in a model system with ferroelectric $\text{PbTiO}_{3}$ as the NC material. $V_{\text{eff}}$ as a function of wavenumber $q$ is shown. Ferroelectric thickness $L_{f}$ could be varied at fixed $n$ to achieve the same effect. The long-wavelength limit is accessed in the case of low carrier density and small thickness.
Energy scales: The low-energy picture described above is appropriate when the adiabatic condition $\hbar\omega_{0}\ll E_{F}$ is satisfied. Fig. 3(a) shows the ratio of 2DES Fermi energy $E_{F}$ to PDT oscillation energy $\hbar\omega_{0}$ (for $4~\text{nm}$ thick $\text{PbTiO}_{3}$) as a function of carrier density. At low density, the adiabatic condition begins to be satisfied ($E_{F}\sim 10\hbar\omega_{0}$) only for linear 2DES with high $v_{F}$. At high density, both linear and parabolic 2DES comfortably satisfy this condition ($E_{F}\sim 100\hbar\omega_{0}$) for typical $v_{F},m^{*}$ values.
The theory developed in this work is most applicable in the low-energy, long-wavelength regime. For this reason, we focus in the remainder of this section on the case of linear 2DES at low density, which according to Fig. 3(a) could be treated as a low-energy system for large enough $v_{F}$ and according to Fig. 2(b) could be modeled reasonably in the long-wavelength approximation (using Eq. (3)).
@@
Figure 3: Critical energy scales and pairing strength. (a) Fermi energy $E_{F}$ of the 2DES, expressed in units of PDT single-mode energy $\hbar\omega_{0}$ for ferroelectric $\text{PbTiO}_{3}$, plotted as a function of Fermi wavenumber $k_{F}$ normalized to the ferroelectric thickness $L_{f}=4~\text{nm}$. The long-wavelength regime described in Fig. 2 is accessed in the limit of low density, shown as encircled region. (b) Pairing strength parameter $\lambda-\mu^{*}$ calculated for a linear 2DES in the long-wavelength approximation as a function of balancing parameter $\zeta$ defined in Eq. (13), shown for two different values of Fermi velocity $v_{F}$. Here, carrier density was fixed at $n=1\times 10^{15}/\text{m}^{2}$. Inset highlights where typical material systems (NC/DE) fall, according to the models used in this work.
@@
From the perspective of real systems, the high density regime and case of parabolic 2DES are just as likely to exhibit interesting physics, but careful treatment of $V_{\text{eff}}(q,\omega)$ in the short-wavelength, high-energy regime is required and beyond the scope of this work.
@@
Pairing strength: To quantify the strength of low-energy electron pairing in the 2DES, we use the parameters $\lambda$ and $\mu^{*}$, adapted from the theory of electron-phonon interaction and superconductivity [colemanbook]. To explain their definition in our context, we begin by decomposing the self-consistently screened interaction $V_{\text{eff}}$ in the following manner, based on the expected slowness of NC-mediated screening relative to direct screening by electrons ($\hbar\omega_{0}\ll E_{F}$):
@@
where $\Pi_{\text{env}}(q)=\Pi_{nc}(q)+\Pi_{d}(q)$ and $V_{\text{el}}(q)=v_{q}/(1-v_{q}\Pi_{\text{el}}(q)$). We call the ‘slow’ term in Eq. (11) $\tilde{V}_{\text{env}}$ and calculate the following dimensionless constants [colemanbook]:
$\begin{aligned} \lambda=-N_{0}\langle\tilde{V}_{\text{env}}\rangle=&-N_{0}\times\frac{\int_{0}^{2\pi}d\theta~\tilde{V}_{\text{env}}(q(\theta))}{2\pi},\\ \mu=N_{0}\langle V_{el}\rangle&=N_{0}\times\frac{\int_{0}^{2\pi}d\theta~V_{el}(q(\theta))}{2\pi},\\ \mu^{*}=&~\frac{\mu}{1+\mu\ln(\frac{E_{F}}{\hbar\omega_{0}})},\end{aligned}$ (12)
where angle brackets denote averaging over Fermi surface (FS), $q(\theta)$ is appropriately defined for the 2D system, and $N_{0}$ is the density of states at $E_{F}$ per spin per unit area. The Coulomb pseudopotential $\mu^{*}$ [morelanderson] represents a renormalization of direct Coulomb repulsion ($\mu$) due to the large bandwidth for electron-electron scattering ($\sim E_{F}$) compared to environment-mediated screening ($\sim\hbar\omega_{0}$). Carrier pairing at accessible temperatures is expected when $\lambda-\mu^{*}>0$ and sufficiently large ($\gtrsim 0.1$).
In the long-wavelength regime, the problem can be fully specified by fixing values for the 2DES $n$ and $v_{F}$, and introducing a balancing parameter $\zeta$ defined as follows:
@@
To satisfy the long-wavelength stability condition in Eq. (5) we assume $\zeta>1$.
@@
Fig. 3(b) shows calculated values of $\lambda-\mu^{*}$ for NC $\text{PbTiO}_{3}$ and linear 2DES ($n=1\times 10^{15}~/\text{m}^{2}$) in the long-wavelength regime. We emphasize two points from this figure. First, the values of $\lambda-\mu^{*}\sim 0.2-0.3$ shown in the inset, calculated using the models described in previous section, are sufficiently large to motivate experimental efforts to observe pairing in such systems. For comparison, in monolayer graphene the value of $\lambda$ arising from other pairing mechanisms such as the electron-phonon interaction is many orders of magnitude smaller [einenkel2011graphene]. Second, balancing the capacitances of NC, DE and 2DES offers a new route to engineer the value of $\lambda-\mu^{*}$. Particularly, as the balancing parameter approaches $\zeta=1$ from above, pairing can be pushed into stronger coupling regime.
Design considerations: From the perspective of NC stabilization, linear and parabolic 2DES systems present qualitative differences. For a linear 2DES, $C_{q}\propto\sqrt{n}/v_{F}$, so the choice of carrier density directly affects the condition for NC stabilization (Eq. (5)). For a parabolic 2DES, $C_{q}$ is independent of carrier density, and scales with effective mass as $C_{q}\propto m^{*}$. This means that effective mass is the critical parameter and that platforms with a larger $m^{*}$ parabolic 2DES will require larger negative $\varepsilon^{nc}$ for NC to be stabilized.
From the perspective of carrier pairing represented by $\lambda-\mu^{*}$, the functional dependence on $n$ and $L_{f}$ is more complicated due to the expressions for FS averaging and retardation effect ($\mu^{*}$). The essential route to maximizing interaction strength is to balance the positive and negative contributions in Eq. (2), and in the long-wavelength regime the problem simplifies considerably, leading to the points made in Fig. 3(b).
V Discussion
@@
In summary, we have presented a framework for Coulomb engineering using negative capacitance materials. Key results include: the idea of Coulomb Engineering using NC (Fig. 1), a stability condition for NC in MF2IM configuration Eq. (5), categorization into regimes based on critical length scales (Fig. 2) and an estimation of energy scales and interaction strength for models of material systems (Fig. 3). We now outline the limitations of our model and discuss directions for future work.
First, the macroscopic Coulomb interaction $V_{\text{eff}}(\vec{q},\omega)$ is generally both wavevector and frequency dependent, and models of this function could be analyzed within a full Green’s function formalism for more quantitative analysis. Our model is best suited for the long-wavelength, static limit $V_{\text{eff}}(q\to 0,\omega\to 0)$, where we have additionally assumed that $V_{\text{eff}}$ is isotropic along the two in-plane directions of 2DES. In the context of engineered phase transitions, detailed knowledge of $V_{\text{eff}}(\vec{q},\omega)$ is required to predict which ordered ground state would be most stable. In the specific case of superconducting ground state, anisotropy of $V_{\text{eff}}$ in the plane of 2DES will affect whether the pairing is expected to be s-wave (as assumed here) or unconventional.
Second, the model for coupling between 2DES and PDT (Kittel model) in Eqs. (6),(7),(8) could be improved in two steps: (i) modeling of optic branch dispersion $\omega_{+}(q)$ (see Appendix C) and coupling $\Omega(q)$ to extend the present result for $\varepsilon^{f}_{z}(q)$ to finite $q$, and (ii) calculation of contribution from all types and branches of excitations in the PDT system [sidorkin2006domain], amounting to complete knowledge of $\varepsilon^{f}_{z}(\vec{q})$ in terms of $\omega_{i}^{L(T)}(\vec{q})$ and $\Omega_{i}^{L(T)}(\vec{q})$ for each branch $i$ and type - longitudinal (L) or transverse (T).
Third, the description of NC in $\text{PbTiO}_{3}$ based on the simple Kittel model may be inadequate when domain structures in experiments exhibit greater complexity (e.g., vortex textures [yadav2019]). Theoretical modeling of the collective excitations in such structures [yang2021condensation] is necessary to account for such details.
Fourth, the models used for 2DES and domain-provided NC are essentially zero-temperature theories. This could be extended by (i) using the Lindhard function at finite temperature [giuliani2008quantum] for $\Pi_{el}$ in Eq. (2) and (ii) incorporating temperature-dependent domain configurations, which would effectively make the NC response function temperature-dependent [pavlenko2021temperature].
Fifth, properties of the bare 2DES itself are expected to deviate significantly from the non-interacting description (used in this work) at low carrier densities. One well-known deviation is negative compressibility [eisenstein1992negative], in which case our assumption of a strictly positive $C_{q}$ from the non-interacting theory may not be a reliable approach. These corrections could be included in the present model by replacing $C_{q}$ with an expression that incorporates interaction effects [giuliani2008quantum].
Finally, we note that the superconducting transition temperature $T_{c}\sim\exp(-1/(\lambda-\mu^{*}))$ so that small errors in estimating the quantity $\lambda-\mu^{*}$ can cause large errors in estimating $T_{c}$. This makes it important to design experiments that can measure small values of $\lambda-\mu^{*}$ directly.
VI Acknowledgements
@@
SD would like to thank Sayeef Salahuddin for alerting him to transient NC and for helpful discussions regarding negative permittivity in ferroelectrics. He would also like to thank Bhaskaran Muralidharan, Kerem Camsari and Zubin Jacob for their feedback on an earlier version of the manuscript. AS and PU gratefully acknowledge helpful discussions with Allan H. MacDonald on anisotropic gate screening.
This work was supported by the Purdue Research Foundation and the Blue Sky Research Program through the Purdue College of Engineering.
@@
Appendix A Anisotropic gate screening effect with Negative Capacitance
@@
We consider a two-dimensional electron system (2DES) encapsulated on one side by a dielectric medium with homogeneous positive static permittivity tensor, $\varepsilon^{d}=\mathrm{diag}(\varepsilon_{\perp}^{d},\varepsilon_{\perp}^{d},\varepsilon_{z}^{d})$, and on the other side by a negative capacitance (NC) medium described by $\varepsilon^{nc}=\mathrm{diag}(\varepsilon_{\perp}^{nc}>0,\varepsilon_{\perp}^{nc}>0,\varepsilon_{z}^{nc}<0)$, where $z$ denotes the stacking direction.
@@
with anisotropy factors $\eta_{d}$ = $\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$ and $\eta_{nc}$ = $\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$. Metal gates are placed at distances $L_{d},L_{nc}$ above and below the $z=0$ plane where 2DES is located.
@@
$\begin{aligned} \varphi(q)\equiv\varphi(q,z=0)&=\alpha^{(nc)}(1-e^{-2\eta_{nc}qL_{nc}})\\ &=\beta^{(d)}(1-e^{-2\eta_{d}qL_{d}}).\end{aligned}$ (A2)
@@
$\begin{aligned} &\alpha^{(nc)}\varepsilon_{z}^{nc}\partial_{z}(e^{\eta_{nc}qz}-e^{-2\eta_{nc}qL_{nc}}e^{-\eta_{nc}qz})\big|_{z=0}\\ &-\beta^{(d)}\varepsilon_{z}^{d}\partial_{z}(e^{-\eta_{d}qz}-e^{-2\eta_{d}qL_{d}}e^{\eta_{d}qz})\big|_{z=0}=\frac{e}{\varepsilon_{0}}.\end{aligned}$ (A3)
Solving for $\alpha^{(nc)}$ and substituting in Eq. (A2), the potential in 2DES is:
@@
Appendix B Stability of Negative Capacitance in MF2IM configuration
@@
Appendix C Domain-provided Negative Capacitance model
In a ferroelectric with periodic domain texture (PDT), the ‘restoring force’ characterizing the response of domain walls to an external electric field can be calculated within the Kittel model of alternating ‘hard’ domains. The equilibrium domain width in this model [catalan2012domain] is:
@@ Appendix C Domain-provided Negative Capacitance model
where $\xi=2\times(1+\varepsilon^{d}_{z}/\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,hf}})$ is a parameter capturing electrostatic boundary conditions, $\delta\approx 1~\text{nm}$ is the domain wall thickness [luk2018prb], and $L_{f}$ is the ferroelectric thickness.
The PDT system has a ‘stiffness’ because there is an energy cost from long-range Coulomb forces when this system is displaced from equilibrium. The oscillation frequency corresponding to this stiffness is calculated using Fourier analysis:
$\begin{aligned} \omega_{\pm}^{2}&(q)=\frac{4P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z}}ML_{f}}\times\\ &\Bigg[\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n-1)^{2}}\bigg)\big(1\pm\cos((2n-1)qd)\big)\\ &-\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n)^{2}}\bigg)(1-\cos(2nqd))\Bigg],\end{aligned}$ (C2)
where the two branches ($\pm$) of excitation for this system are analogous to the acoustic and optic branches in crystals with a two-atom basis. In this work we restrict attention to the ++ (‘optic’) branch in the long-wavelength limit, $q\to 0$, so $\omega_{0}$ in the main text is just $\omega_{+}(q=0)$. Within this model, the calculated negative permittivity $\varepsilon^{f}_{z}$ has a value of about 60-60 for $L_{f}=4~\text{nm}$.

### v2

@@ Abstract
[…]
@@
Figure 3: Critical energy scales and pairing strength. (a) Fermi energy $E_{F}$ of the 2DES, expressed in units of PDT single-mode energy $\hbar\omega_{0}$ for ferroelectric $\text{PbTiO}_{3}$, plotted as a function of Fermi wavenumber $k_{F}$ normalized to the ferroelectric thickness $L_{f}$. The long-wavelength regime described in Fig. 2 is accessed in the limit of low density, shown as circled region. (b) Pairing strength parameter $\lambda-\mu^{*}$ calculated for a linear 2DES in the long-wavelength approximation as a function of balancing parameter $\zeta$ defined in Eq. (11), shown for two typical values of Fermi velocity $v_{F}$. Here, carrier density was fixed at $n=1\times 10^{15}/\text{m}^{2}$. Inset highlights where typical material systems (NC/DE) fall, according to the models used in this work.
In frequency space, the negative permittivity associated with the PDT response [Eq. (5)] persists below the characteristic PDT oscillation frequency $\omega_{0}$ [39]. Consequently, the Coulomb interaction mediated by the NC ferroelectric remains attractive for $\omega<\omega_{0}$. When the Fermi energy satisfies $E_{F}\gg\hbar\omega_{0}$, this separation of energy scales allows the retarded attractive interaction to partially overcome the instantaneous Coulomb repulsion, analogous to phonon-mediated pairing. Motivated by this observation, we next estimate the pairing strength within our model platform.
@@
Fig. 3(a) shows the ratio of 2DES Fermi energy $E_{F}$ to PDT oscillation energy $\hbar\omega_{0}$ (for $4~\text{nm}$ thick $\text{PbTiO}_{3}$) as a function of carrier density. At low density, this condition begins to be satisfied ($E_{F}\sim 10\hbar\omega_{0}$) only for linear 2DES with high $v_{F}$. At high density, both linear and parabolic 2DES comfortably satisfy this condition ($E_{F}\sim 100\hbar\omega_{0}$) for typical $v_{F},m^{*}$ values. The theory developed in this work is most applicable in the low-energy, long-wavelength regime. For this reason, we focus in the remainder of this section on the case of linear 2DES at low density, which we model in the long-wavelength approximation using Eq. (3). From the perspective of real systems, the high density regime and case of parabolic 2DES are just as likely to exhibit interesting physics, but careful treatment of $V_{\text{eff}}(q,\omega)$ in the short-wavelength, high-energy regime is required and beyond the scope of this work.
@@
To quantify the strength of low-energy electron pairing in the 2DES, we use parameters $\lambda$ and $\mu^{*}$, adapted from the theory of electron-phonon interaction and superconductivity [8]. To explain their definition in our context, we begin by decomposing the self-consistently screened interaction $V_{\text{eff}}$ in the following manner, based on the expected slowness of NC-mediated screening relative to direct screening by electrons ($\hbar\omega_{0}\ll E_{F}$):
@@
where $\Pi_{\text{env}}(q)=\Pi_{nc}(q)+\Pi_{d}(q)$ and $V_{\text{el}}(q)=v_{q}/(1-v_{q}\Pi_{\text{el}}(q)$). We call the ‘slow’ term in Eq. (9) $\tilde{V}_{\text{env}}$ and calculate the following dimensionless constants [8]:
$\lambda=-N_{0}\langle\tilde{V}_{\text{env}}\rangle\hskip 1.28032pt;\hskip 1.28032pt\mu=N_{0}\langle V_{el}\rangle\hskip 1.28032pt;\hskip 1.28032pt\mu^{*}=\frac{\mu}{1+\mu\ln(\frac{E_{F}}{\hbar\omega_{0}})},$ (10)
where angle brackets denote averaging over Fermi surface (FS), $q(\theta)$ is appropriately defined for the 2D system, and $N_{0}$ is the density of states at $E_{F}$ per spin per unit area. The Coulomb pseudopotential $\mu^{*}$ [41] represents a renormalization of direct Coulomb repulsion ($\mu$) due to the large bandwidth for electronic screening ($\sim E_{F}$) compared to environment-mediated screening ($\sim\hbar\omega_{0}$). Carrier pairing at accessible temperatures is expected when $\lambda-\mu^{*}>0$ and sufficiently large ($\gtrsim 0.1$).
In the long-wavelength limit, the problem can be fully specified by fixing values for the 2DES $n$ and $v_{F}$, and introducing a balancing parameter $\zeta$ defined as follows:
@@
To satisfy the thermodynamic stability condition in Eq. (4) we assume $\zeta>1$. Fig. 3(b) shows calculated values of $\lambda-\mu^{*}$ for NC $\text{PbTiO}_{3}$ and linear 2DES ($n=1\times 10^{15}~/\text{m}^{2}$) in the long-wavelength limit. We emphasize two points from this figure. First, estimates for the model platform introduced above yield $\lambda-\mu^{*}\gtrsim 0.1$ (inset). Although these estimates rely on simplifying assumptions, discussed further in the next section, they suggest that pairing strengths of experimentally relevant magnitude may be achievable. For comparison, in monolayer graphene the value of $\lambda$ arising from other pairing mechanisms such as the electron-phonon interaction is many orders of magnitude smaller [12]. Second, in contrast to phonon-mediated pairing where the coupling strength is typically fixed by material properties, balancing the capacitances of NC, DE and 2DES offers a new route to engineer the value of $\lambda-\mu^{*}$. Particularly, as the balancing parameter approaches $\zeta=1$ from above, pairing can be pushed into strong coupling regimes [^1] We note that the assumptions of linear response and the existence of a small parameter $\lambda\hbar\omega_{0}/E_{F}$, which underlie the present pairing estimates, may break down as $\zeta\rightarrow 1$. This could ultimately limit the achievable pairing strength..
@@
Discussion.— In summary, we have introduced a new method for Coulomb engineering using negative-capacitance materials. We show that incorporating NC media enables the effective electron–electron interaction in two-dimensional systems to be tuned and even reversed in sign while maintaining thermodynamic stability (Fig. 1). We additionally present a stability condition for NC in the MF2IM configuration [Eq. (4)], categorization of Coulomb engineering into regimes based on critical length scales(Fig. 2), and estimates of interaction strengths for realistic material platforms (Fig. 3). We now outline the limitations of our analysis and discuss directions for future work.
A more quantitative description of $V_{\mathrm{eff}}(\vec{q},\omega)$ requires extending the present treatment beyond the long-wavelength and low-frequency limits considered here. For the ferroelectric NC medium, this includes: (i) incorporating dispersion of the optic branch $\omega_{+}(q)$ and coupling $\Omega(q)$ to extend the present results for $\varepsilon_{z}^{f}(q=0)$ to finite $q$ (see Appendix C); (ii) including contributions from all excitation branches in the PDT system [55]; (iii) accounting for anisotropy introduced by domain structures, which may lead to angular dependence of $\epsilon_{z}^{f}(\vec{q})$ and influence pairing symmetry; (iv) extending the description of NC beyond the Kittel model [62], particularly given experimentally observed vortex-like structures [61]; and (v) incorporating the temperature dependence of the NC phase and its dielectric response [45].
For the 2DES, further refinements include incorporating finite-thickness corrections through $q$-dependent form factors [1] and extending $\Pi_{\rm el}$ beyond the static, zero-temperature limit by using finite-temperature dynamical screening approaches [16]. This extension incorporates the plasmon, which is important even in the low-frequency limit due to its gapless nature in two dimensions. Hybridization [23] of this mode with excitations of the NC medium can modify $V_{\mathrm{eff}}(\vec{q},\omega)$ and conclusions about the resulting ground states.
Finally, the estimates of pairing strength using Eq. (10) can be improved by analyzing the full momentum- and frequency-dependent interaction $V(\vec{q},\omega)$ within a Green’s function framework [14, 12]. When pair-breaking processes limit superconductivity, then the transition temperature scales as $T_{c}\sim\exp(-1/(\lambda-\mu^{*}))$, so that small errors in estimating the quantity $\lambda-\mu^{*}$ can cause large errors in estimating $T_{c}$. This makes it important to design experiments capable of measuring small values of $\lambda-\mu^{*}$ directly. Moreover, in two dimensions $T_{c}$ may instead be limited by phase fluctuations, highlighting the need to estimate the superfluid stiffness [50] in future work.
@@
Acknowledgements.—SD would like to thank Sayeef Salahuddin for alerting him to transient NC and for helpful discussions regarding negative permittivity in ferroelectrics. He would also like to thank Bhaskaran Muralidharan, Kerem Camsari and Zubin Jacob for their feedback on an earlier version of the manuscript. AS and PU gratefully acknowledge helpful discussions with Allan H. MacDonald on anisotropic gate screening. This work was supported by the Purdue Research Foundation and the Blue Sky Research Program through the Purdue College of Engineering.
Appendix A APPENDIX
@@
Appendix B A. Anisotropic gate screening effect with Negative Capacitance
@@
We consider a two-dimensional electron system (2DES) encapsulated on one side by a dielectric medium with homogeneous positive static permittivity tensor, $\varepsilon^{d}=\mathrm{diag}(\varepsilon_{\perp}^{d},\varepsilon_{\perp}^{d},\varepsilon_{z}^{d})$, and on the other side by a negative-capacitance (NC) medium described by $\varepsilon^{nc}=\mathrm{diag}(\varepsilon_{\perp}^{nc}>0,\varepsilon_{\perp}^{nc}>0,\varepsilon_{z}^{nc}<0)$, where $z$ denotes the stacking direction.
@@
with anisotropy factors $\eta_{d}$ = $\sqrt{\varepsilon^{d}_{\perp}/\varepsilon^{d}_{z}}$ and $\eta_{nc}$ = $\sqrt{\varepsilon^{nc}_{\perp}/\varepsilon^{nc}_{z}}$. Metal gates are placed at distances $L_{d},L_{nc}$ above and below the $z=0$ plane where the 2DES is located.
@@
$\begin{aligned} \varphi(q)\equiv\varphi(q,z=0)&=\alpha^{(nc)}(1-e^{-2\eta_{nc}qL_{nc}})\\ &=\beta^{(d)}(1-e^{-2\eta_{d}qL_{d}}).\end{aligned}$ (13)
@@
$\begin{aligned} \alpha^{(nc)}&\varepsilon_{z}^{nc}\partial_{z}(e^{\eta_{nc}qz}-e^{-2\eta_{nc}qL_{nc}}e^{-\eta_{nc}qz})\big|_{z=0}\\ -&\beta^{(d)}\varepsilon_{z}^{d}\partial_{z}(e^{-\eta_{d}qz}-e^{-2\eta_{d}qL_{d}}e^{\eta_{d}qz})\big|_{z=0}=e/\varepsilon_{0}.\end{aligned}$ (14)
Solving for $\alpha^{(nc)}$ and substituting in Eq. (13), the potential in 2DES is:
@@
Appendix C B. Stability of Negative Capacitance in MF2IM configuration
@@
An important consideration neglected in this work is the modification of $C_{q}$ (and $\Pi_{el}$ in main text) by ‘direct’ electron-electron interactions (the screening channel constituted by identical electrons within the 2DES). The strength of this interaction relative to kinetic energy is often expressed in terms of the dimensionless parameter $r_{s}$. For a 2DES with spin and valley degeneracies $g_{s}$ and $g_{v}$ and background dielectric constant $\kappa$, we have the following expressions for different cases of dispersion relation [9]:
$r_{s}=\begin{cases}\displaystyle\frac{e^{2}}{4\pi\varepsilon_{0}\kappa\hbar v_{F}}\,\frac{\sqrt{g_{s}g_{v}}}{2},&\text{Linear},\\[6.0pt] \displaystyle\frac{m^{*}e^{2}}{8\pi\varepsilon_{0}\kappa\hbar^{2}}\,\frac{g_{s}g_{v}}{\sqrt{\pi n}},&\text{Parabolic}.\end{cases}$ (20)
For a 2DES at large enough $r_{s}$, the modified $C_{q}$ could itself be negative due to exchange-correlation corrections [13, 16]. In that case, Eqs. (17) and (18) remain valid, but Eq. (19) becomes only a necessary (not sufficient) condition, and Eq. (4) of the main text must be revised. A more detailed analysis indicates that, within the model used in this section, an MF2IM structure with both $C_{nc}<0$ and $C_{q}<0$ would actually be unstable, in the sense that no simple necessary and sufficient stability condition can be formulated. It is nevertheless conceivable that an appropriately designed stabilizing circuit could enable access to this regime, should it prove to be experimentally relevant.
Appendix D C. Domain-provided Negative Capacitance model
In a ferroelectric with periodic domain texture (PDT), the ‘restoring force’ characterizing the response of domain walls to an external electric field can be calculated within the Kittel model [33] of alternating ‘hard’ domains. The equilibrium domain width in this model [7] is:
@@
where $\xi=2\times(1+\varepsilon^{d}_{z}/\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z,hf}})$ is a parameter capturing electrostatic boundary conditions, $\delta\approx 1~\text{nm}$ is the domain wall thickness [39], and $L_{f}$ is the ferroelectric thickness.
The PDT system has a ‘stiffness’ because there is an energy cost from long-range Coulomb forces when this system is displaced from equilibrium. The oscillation frequency corresponding to this stiffness is calculated in the Kittel model [55] using Fourier analysis:
$\begin{aligned} \omega_{\pm}^{2}&(q)=\frac{4P_{s}^{2}}{\pi\varepsilon_{0}\sqrt{\varepsilon^{f}_{\perp}\varepsilon^{f}_{z}}ML_{f}}\times\\ &\Bigg[\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n-1)^{2}}\bigg)\big(1\pm\cos((2n-1)qd)\big)\\ &-\sum_{n=1}^{\infty}\textrm{ln}\bigg(1+\frac{\varepsilon^{f}_{\perp}}{\varepsilon^{f}_{z}}\frac{L_{f}^{2}}{d^{2}}\frac{1}{(2n)^{2}}\bigg)(1-\cos(2nqd))\Bigg],\end{aligned}$ (22)
where the two branches ($\pm$) of excitation for this system are analogous to the acoustic and optic branches in crystals with a two-atom basis. In this work we restrict attention to the ++ (‘optic’) branch in the long-wavelength limit, $q\to 0$. We calculate $\omega_{0}$ in the main text as $\omega_{+}(q=0)$ and assume electric dipole coupling of the form $P_{z}\times E_{z}$ [55, 39]:
Parameters for ferroelectric $\text{PbTiO}_{3}$ are chosen as follows [39] : $P_{s}=0.65~\text{Cm}^{-2}$ is the spontaneous polarization, and $M$ is the domain wall mass per unit area estimated using interpolation formula $M=1.3\sqrt{L_{f}[\text{nm}]}\times 10^{-9}~\text{kgm}^{-2}$ [39, 64]. $\varepsilon^{f}_{z,hf}=100$ is the high-frequency dielectric constant capturing the background polarizability of $\text{PbTiO}_{3}$ arising from other polarization mechanisms in the material, and $\varepsilon^{f}_{\perp}=30$.
Within this model, the calculated negative permittivity $\varepsilon^{f}_{z}$ has a value of about 60-60 for $L_{f}=4~\text{nm}$.